# RCooper: A Real-world Large-scale Dataset for Roadside Cooperative Perception

### v1

@@ 1 Introduction
Previous roadside perception [43, 41, 42, 14] mainly concentrates on the perception from an independent roadside view due to the accessible dataset [44, 45]. However, they can not realize a comprehensive understanding of a traffic area. Single-infrastructure roadside perception is constrained by the installation perspective, leading to the limited sensing range and blind spots, which can be handled via cross-infrastructure cooperation. Observation from various views can extend the sensing range, reduce blind spots, and further enhance the understanding of the same instance. Towards practical applications, Roadside Cooperative Perception (RCooper) is expected to achieve area-coverage roadside perception for restricted traffic areas, which is shown in Fig. 1. The capabilities of RCooper, including extending the sensing range and reducing blind spots, are illustrated in Fig. 2.
Technically, three challenges can be derived from RCooper: 1) Data heterogeneity. Considering the construction cost, various types of sensors (multiline LiDAR, MEMS LiDAR, and camera) are employed in practice, leading to prominent data heterogeneity for cooperative perception [4, 16]. 2) Cooperative representation merits further enhancement. Most existing cooperative perception approaches, including well-explored vehicle-vehicle (V2V) and vehicle-infrastructure (V2I) cooperation, are designed for vehicle-centric cooperative tasks [28, 9, 31, 50, 24, 34, 46]. However, to our knowledge, the roadside cooperation approaches have not been investigated before. The inherent characteristics of roadside sensors (like roll, pitch angle, height) make roadside cooperative representation a different playground compared with vehicle-centric cooperation [44, 42, 43]. 3) Perception performance needs improvements. How to achieve high-quality downstream perception tasks based on roadside cooperative representation, e.g., detection, tracking, counting, and monitoring, needs further investigation. For example, tracking by unstable detection results in complex intersection scenes is still challenging. It is necessary to delve into these challenges, but the lack of datasets hinders the exploration of the Garden of Eden.
@@ 2 Related work
Dataset type
Source Dataset Year Coop Mode RGBs LiDARs Catagories Det task Trk task
Roadside
real BoxCars [32] 2018 None 116k - 12 - -
@@ 3.1 Data Acquisition
  Sensors
Details
  Camera RGB
42Hz, $1920\times 1200$
LiDAR 80-beams
10Hz, $360^{\circ}$ horizontal FOV, $-25^{\circ}$ to $15^{\circ}$ vertical FOV, 1m to 230m capture range, ±$\pm 3cm$ error
32-beams
10Hz, $360^{\circ}$ horizontal FOV, $90^{\circ}$ vertical FOV, 0.1m to 30m capture range, ±$\pm 3cm$ error
MEMS
10Hz, ±$\pm 60^{\circ}$ horizontal FOV, ±$\pm 12.5^{\circ}$ vertical FOV, 0.5m to 200m capture range, ±$\pm 3cm$ error
@@ 5.2 Roadside Cooperative Detection Results
  Method AMOTA normal-$\uparrow$ AMOTP normal-$\uparrow$ sAMOTA normal-$\uparrow$ MOTA normal-$\uparrow$ MT normal-$\uparrow$ ML normal-$\downarrow$
@@
  Method AMOTA normal-$\uparrow$ AMOTP normal-$\uparrow$ sAMOTA normal-$\uparrow$ MOTA normal-$\uparrow$ MT normal-$\uparrow$ ML normal-$\downarrow$

### v2

@@ 1 Introduction
Previous roadside perception [44, 42, 43, 13, 21, 53] mainly concentrates on the perception from an independent roadside view due to the accessible dataset [45, 46]. However, they can not realize a comprehensive understanding of a traffic area. Single-infrastructure roadside perception is constrained by the installation perspective, leading to the limited sensing range and blind spots, which can be handled via cross-infrastructure cooperation. Observation from various views can extend the sensing range, reduce blind spots, and further enhance the understanding of the same instance. Towards practical applications, Roadside Cooperative Perception (RCooper) is expected to achieve area-coverage roadside perception for restricted traffic areas, which is shown in Fig. 1. The capabilities of RCooper, including extending the sensing range and reducing blind spots, are illustrated in Fig. 2.
Technically, three challenges can be derived from RCooper: 1) Data heterogeneity. Considering the construction cost, various types of sensors (multiline LiDAR, MEMS LiDAR, and camera) are employed in practice, leading to prominent data heterogeneity for cooperative perception [4, 15]. 2) Cooperative representation merits further enhancement. Most existing cooperative perception approaches, including well-explored vehicle-vehicle (V2V) and vehicle-infrastructure (V2I) cooperation, are designed for vehicle-centric cooperative tasks [28, 39, 31, 51, 24, 34, 47]. However, to our knowledge, the roadside cooperation approaches have not been investigated before. The inherent characteristics of roadside sensors (like roll, pitch angle, height) make roadside cooperative representation a different playground compared with vehicle-centric cooperation [45, 43, 44, 52]. 3) Perception performance needs improvements. How to achieve high-quality downstream perception tasks based on roadside cooperative representation, e.g., detection, tracking, counting, and monitoring, needs further investigation. For example, tracking by unstable detection results in complex intersection scenes is still challenging. It is necessary to delve into these challenges, but the lack of datasets hinders the exploration of the Garden of Eden.
@@ 2 Related work
  Dataset type Source Dataset Year Coop Mode RGBs LiDARs Catagories Det task Trk task
  Roadside real BoxCars [32] 2018 None 116k - 12 - -
@@ 3.1 Data Acquisition
                      Sensors Details
  Camera RGB 42Hz, $1920\times 1200$
LiDAR 80-beams 10Hz, $360^{\circ}$ horizontal FOV, $-25^{\circ}$ to $15^{\circ}$ vertical FOV, 1m to 230m capture range, ±$\pm 3cm$ error
32-beams 10Hz, $360^{\circ}$ horizontal FOV, $90^{\circ}$ vertical FOV, 0.1m to 30m capture range, ±$\pm 3cm$ error
MEMS 10Hz, ±$\pm 60^{\circ}$ horizontal FOV, ±$\pm 12.5^{\circ}$ vertical FOV, 0.5m to 200m capture range, ±$\pm 3cm$ error
@@ 5.2 Roadside Cooperative Detection Results
  Method AMOTA $\uparrow$ AMOTP $\uparrow$ sAMOTA $\uparrow$ MOTA $\uparrow$ MT $\uparrow$ ML $\downarrow$
@@
  Method AMOTA $\uparrow$ AMOTP $\uparrow$ sAMOTA $\uparrow$ MOTA $\uparrow$ MT $\uparrow$ ML $\downarrow$
@@ 6 Conclusion
Supplementary Material
A Appendix
This supplementary document is organized as follows:
  • Additional information and visualization of RCooper is shown in Appendix B.
  • The visualization of roadside cooperative perception results is supplemented in Appendix C, offering an illustrative overview of the performance evaluation of our established benchmark.
  • Some preliminary camera-based experiment results is shown in Appendix D.
B Dataset Visualization
More visualizations of the proposed RCooper dataset are shown in Fig 7. These visualizations serve as a valuable supplementary to the discussion in Section 3.1, where two representative traffic scenes (i.e., intersection and corridor) are presented. Each figure within this context comprises an aggregated 3D LiDAR point cloud image and four camera images captured from distinct views. For the corridor scenes, the point clouds generated by two different LiDAR sensors are depicted in red and green, respectively. The infrastructure-side cooperation can effectively extend the sensing range to cover the whole corridor scene. For the intersection scenes, the point clouds generated by multiline LiDARs and MEMES LiDARs are denoted in red and green, respectively. The observations from multiple views serve to mitigate the challenges posed by occlusions in intricate intersection scenarios.
Figure 7: Visualization Examples of the RCooper Dataset.
C Perception Results Visualization
More visualizations of benchmark results are shown in this section.
Typical perception benchmark results for corridor scenes (referred to as CoBEVT results) and intersection scenes (referred to as late fusion results) are illustrated in Fig 8 and Fig 9, respectively. Within each figure, alongside the LiDAR points and a camera image, we have delineated the predicted objects with red outlines and the ground truth objects with yellow outlines. For the corridor scenes (in Fig 8) and uncrowded intersection scenes (in Fig 9. a, b, c and d), the perception process proves to be highly effective. For the crowded intersection scenes (in Fig 9. e, f, g and h), the presence of occlusions and the inherent complexity of the scenarios introduce false negatives and false positives into the perception process. These challenges merit further investigation in future research endeavors.
Figure 8: Visualization Examples of the Perception Results for Corridor Scenes. Red and yellow bounding boxes represent the prediction and ground truth, respectively.
Figure 9: Visualization Examples of the Perception Results for Intersection Scenes. Red and yellow bounding boxes represent the prediction and ground truth, respectively.
D Preliminary Camera-based Experiment Results
We provide some preliminary camera-based experiment results in this section. A single-view camera suffers from the inherent blind spot, which can be theoretically mitigated by roadside cooperative perception. Camera-based late-fusion results in both corridor and intersection scenes are reported in Tab. 7. It’s worth noting that the data annotation of our datasets is based on point cloud data, and meanwhile the sensing range of camera group is smaller than that of LiDAR group, which means that the label may need filtering on demand considering camera group’s view.
  Scenario 3D@0.3 3D@0.5 BEV@0.3 BEV@0.5
  Corridor 26.99 3.53 38.70 16.15
Intersection 11.32 2.41 12.19 7.89
Table 7: Camera-based cooperative detection AP metric (%).